OpenStudy (anonymous):

Suppose you walk straight from the jaguars to the tigers and then straight to the otters. What is the total distance you walked, to the nearest tenth of a unit?

OpenStudy (anonymous):

is this all the info provided?

OpenStudy (anonymous):

this aswell

OpenStudy (anonymous):

have you learned about the distance formula?

OpenStudy (anonymous):

yes i have, but awhile ago, can you refresh my memory?

OpenStudy (anonymous):

suppose you've two points (x1,y1),(x2,y2) then distance between them d=sqrt[(x1-x2)^2+(y1-y2)^2]

OpenStudy (anonymous):

so plug in the x2 and y2 as the tigers data points and x1 and y1 for the jags

OpenStudy (anonymous):

this is where i get confused

OpenStudy (anonymous):

where the jag dot is (what its coordinates are?)

OpenStudy (anonymous):

6, -5 ?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

yeah do you know where the tigers is

OpenStudy (anonymous):

(4, -1)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what did i say before about jag and tigers points (x1 y1 x2 y2) so now you know what each one equals..then just plug it in..

OpenStudy (anonymous):

d= 4.4?

OpenStudy (anonymous):

\[2\sqrt{5} or \sqrt{20}\]

OpenStudy (anonymous):

close enough though

OpenStudy (anonymous):

but it isnt over yet lol

OpenStudy (anonymous):

you have to do the same for tiger to otter

OpenStudy (anonymous):

can u handle that by urself?

OpenStudy (anonymous):

d= \[\sqrt{89}\] ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay then i add those, correct or no?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Thank you! [:

OpenStudy (anonymous):

u r actually gonna have to plug it into a calc or put into decimal form because 89 is prime

OpenStudy (anonymous):

well, with the A,B,C,D choices I have, one of them makes the most sense to me, "13.9"

OpenStudy (anonymous):

yeah seems closest

OpenStudy (anonymous):

Okay, well thanks again (: